### v1

# RESSA: Repair Sparse Vision-Language Models via Sparse Cross-Modality Adaptation

Abstract
Vision-Language Models (VLMs), integrating diverse information from multiple modalities, have shown remarkable success across various tasks. However, deploying VLMs, comprising large-scale vision and language models poses challenges in resource-constrained scenarios. While pruning followed by finetuning offers a potential solution to maintain performance with smaller model sizes, its application to VLMs remains relatively unexplored, presenting two main questions: how to distribute sparsity across different modality-specific models, and how to repair the performance of pruned sparse VLMs. To answer the first question, we conducted preliminary studies on VLM pruning and found that pruning vision models and language models with the same sparsity ratios contribute to nearly optimal performance. For the second question, unlike finetuning unimodal sparse models, sparse VLMs involve cross-modality interactions, requiring specialized techniques for post-pruning performance repair. Moreover, while parameter-efficient LoRA finetuning has been proposed to repair the performance of sparse models, a significant challenge of weights merging arises due to the incompatibility of dense LoRA modules with sparse models that destroy the sparsity of pruned models. To tackle these challenges, we propose to Repair Sparse Vision-Language Models via Sparse Cross-modality Adaptation (RESSA). RESSA utilizes cross-modality finetuning to enhance task-specific performance and facilitate knowledge distillation from original dense models. Additionally, we introduce SparseLoRA, which applies sparsity directly to LoRA weights, enabling seamless integration with sparse models. Our experimental results validate the effectiveness of RESSA, showcasing significant enhancements, such as an 11.3% improvement under 2:4 sparsity and a remarkable 47.6% enhancement under unstructured 70% sparsity. Code will be released at: https://github.com/Shwai-He/RESSA.
1 Introduction
Scaling deep learning models has demonstrated promising performance across various tasks in both vision and language domains Brown et al. (2020); Jiang et al. (2024); Zhu et al. (2023). Vision-Language Models (VLMs) Radford et al. (2021); Li et al. (2022); Liu et al. (2023), which leverage powerful vision and language models, have recently garnered significant attention in research OpenAI et al. (2024); Liu et al. (2024), showcasing their cross-modality capabilities. However, the ever-increasing size of these models comes with substantial computational and memory costs, limiting their practical applicability in resource-constrained environments. Model pruning followed by finetuning Dai et al. (2018); Fang et al. (2023); Tanaka et al. (2020), which reduces model size while preserving performance, holds promise for improving the real-world deployment of VLMs.
While pruning followed by finetuning has significantly improved the efficiency of vision models Frankle & Carbin (2019); Kusupati et al. (2020); Lee et al. (2019) and language models Chen et al. (2020); Sun et al. (2023a); Frantar & Alistarh (2023), the realm of Vision-Language Models (VLMs) remains relatively unexplored in terms of model pruning, prompting following questions: how to distribute sparsity ratios between different modality-specific models and how to repair the performance of prune sparse VLMs.
For the first question, we conducted empirical studies focusing on pruning modality-specific models. Specifically, we maintained a constant summation of sparsity ratios for language models and vision models and then experimented with various combinations of sparsity ratios. Surprisingly, we discovered that pruning the vision model and language model with the same sparsity ratios contributes to nearly optimal performance. However, even so, pruning both vision models and language models simultaneously would significantly degrade performance, especially for structured sparsity patterns (e.g., N:M sparsity Zhang et al. (2022); andYukun Ma et al. (2021)). While this degradation can be mitigated by only pruning language models, there still exists a significant performance gap compared to the full model, demonstrating the importance of post-pruning repairing.
Unlike finetuning unimodal sparse models, sparse VLMs involve cross-modality interactions, requiring specialized techniques for post-pruning performance repair. Moreover, while parameter-efficient LoRA finetuning has been proposed to repair the performance of sparse models, a significant challenge arises due to the incompatibility of dense LoRA modules with sparse models: merging LoRA modules with sparse models would destroy the sparse pattern while maintaining LoRA modules would cause extra latency and slow down the inference speed.
In response to the above challenges, we propose Repair Sparse Vision-Language Models via Sparse Cross-Modality Adaptation (RESSA). Specifically, we define two repairing objectives: enhancing task-specific performance and minimizing divergence from the original dense model. The former objective guides the model to improve performance, while the latter facilitates knowledge transfer from dense models to pruned models. To address the incompatibility issue of LoRA, we introduce SparseLoRA finetuning, which utilizes binary masks on LoRA weights, allowing seamless integration with pruned weights. SparseLoRA is incorporated into both the vision and language model to ensure cross-modality finetuning.
Extensive experiments showcase the effectiveness of our proposed RESSA in repairing the performance of pruned sparse VLMs. For instance, as illustrated in Figure 1, RESSA boosts the performance by 13.1% for InstructBLIP-Vicuana-7B with 2:4 sparsity. In summary, our contributions are threefold:
We empirically study the modality-specific sparsity distributions and systematically demonstrate how sparsity affects the performance of VLMs.
Following the empirical studies, we propose RESSA that repairs the pruned models via cross-modality adaptation and SparseLoRA finetuning.
Extensive experiments demonstrate the effectiveness and universality of RESSA across various VLMs and tasks.
2 Related Work
Vision-Language Models. Vision-language models, among the most sophisticated multi-modal architectures, have demonstrated outstanding performance across various cross-modality tasks, including image captions Sharma et al. (2018), image retrieval Plummer et al. (2015), visual QA Kim et al. (2016), and image/video generation Zhou et al. (2021); Singer et al. (2022). These models typically freeze the pretrained vision and language components, only fine-tuning a small, learnable interface (e.g., Qformer in BLIP-2 Li et al. (2023)) to facilitate inter-modality interactions Yin et al. (2023); Li et al. (2023), thus avoiding high training costs and potential catastrophic forgetting Goodfellow et al. (2014).
Model Pruning for Large Language Models. While large vision and language models have shown promising advancements, their massive parameter sizes present challenges for practical deployment Ma et al. (2023); Wang et al. (2020). To mitigate this, model pruning techniques have been introduced to remove redundant weights or structures Han et al. (2016); Alvarez & Salzmann (2016). The primary aim of model pruning is to minimize the disparity between models before and after pruning Liu et al. (2021); He & Xiao (2024); Frantar & Alistarh (2023). Various metrics, such as magnitude, gradient Yi-Lin Sung (2024), and activation Sun et al. (2023a), have been proposed to identify unimportant weights. However, pruning without finetuning often leads to a performance drop. Zhang et al. (2024) utilize reconstruction errors-based metrics to update the weights. Other than the disparity between sparse models and dense models, our method also considers the task-specific objective of repairing sparse models and knowledge distillation from the original full models.
3 Preliminary Study
Vision-Language Models (VLMs) consist of modality-specific foundation models, namely visual and language models, as well as a cross-modality interface (e.g., QFormer Li et al. (2023)) that aligns models from different modalities. Following Yi-Lin Sung (2024), we focus on pruning the vision and language models while keeping the Q-Former intact, as it is sufficiently lightweight. Parameters are not evenly distributed across the different modality-specific models; for instance, visual models are often considerably smaller than their corresponding language models Li et al. (2023); Dai et al. (2023); Liu et al. (2023); Yang et al. (2022). In this case, we pose two questions: (1) how to distribute sparsity ratios between modality-specific models, and (2) how do different sparsity ratios affect the performance of VLMs?
For the first question, we first try various sparsity ratio combinations between visual models and language models. Specifically, we fix the summation of and and then adjust their distributions accordingly. We use Wanda Sun et al. (2023a) as the default pruning method because it ensures relatively high performance and efficiency. Based on Figure 2, we found: (1) VLMs would collapse when the language models are under high sparsity ratios (i.e., ), whereas sparsity imposed on visual models has a comparatively lower impact on performance; (2) When constrained by the summation of sparsity (), pruning the modality-specific models with equal sparsity ratios leads to optimal performance.
For the second question, we initially prune VLMs with different unstructured sparsity ratios using the following strategies: pruning language models and visual models with the same sparsity ratios (“V + L”), pruning visual models only (“Vision”), and pruning language models only (“Language”). According to Figure 3, when sparsity ratios exceed 50%, all settings experience a significant performance drop, although VLMs pruned by a single modality model maintain relatively high performance.
Similarly, in Figure 3 when employing structured N:M sparsity andYukun Ma et al. (2021); Zhang et al. (2022) (i.e., in each contiguous block of values, values must be zero), all models encounter significant performance degradation and even collapse (2:4 for pruning both vision models and language models). This situation prompts us to contemplate the question of how to repair the performance degradation caused by pruning for VLMs.
4 Methodology
In this section, we will develop a pruning pipeline, RESSA, that repairs the pruned models via cross modality adaptation and SparseLoRA finetuning, with the illustration in Figure 4.
4.1 Pruning with Few Samples
Model pruning identifies less important weights using predefined metrics Han et al. (2016), typically measuring the reconstruction errors Sun et al. (2023a); Frantar & Alistarh (2023); Zhang et al. (2024) between models before and after pruning, such as magnitude, gradient, and activation. Calculating gradients or activation requires a small calibration dataset with few samples. With predefined metric and the calibration dataset, the weights of a model is scored as follows:
where denote the weights of the model while represent the importance scores for . Given the sparse ratios , binary masks are utilized to locate the pruned weights and update the weights as follows:
where denotes the pruned weights, while represents the threshold ( percentile of ) and all weights with scores lower than will be removed. While the pruning metrics aim to minimize reconstruction errors Sun et al. (2023a); Zhang et al. (2024) or maintain performance Yi-Lin Sung (2024), model pruning often results in a significant performance drop and therefore needs to be recovered.
4.2 Sparse LoRA finetuning
VLMs, which incorporate both vision models and language models, are often too large to be fine-tuned through full-model fine-tuning techniques Li et al. (2023); Dai et al. (2023). Instead, parameter-efficient fine-tuning techniques Houlsby et al. (2019); Mangrulkar et al. (2022); Hu et al. (2022) are employed to reduce the number of trainable parameters while maintaining comparable performance. Among these techniques, LoRA Hu et al. (2022) stands out as one of the most widely used approaches. since it not only efficiently utilizes parameters but also allows for seamless integration with the original weights, thus avoiding potential latency during inference Dery et al. (2024); Rücklé et al. (2021).
Traditional LoRA fine-tuning involves freezing the parameters of the pretrained model and injecting trainable rank decomposition matrices into each layer that requires fine-tuning. LoRA modules involves two small low-rank trainable weights and , which can be merged with after finetuning:
However, as shown in Figure 5, the sparse pattern of pruned models would collapse after merging Dery et al. (2024); He et al. (2023). Given that are dense weights and are sparse weights, the element-wise operation would destroy the sparse patterns. Additionally, without merging, the injected LoRA modules would increase latency and slow down inference speed Mundra et al. (2023); Rücklé et al. (2021); Dery et al. (2024). Inspired by He et al. (2022), we propose employing masks on to preserve the sparse pattern:
In such a case, corresponding to pruned positions are masked and cannot be updated via gradient-backpropagation. Consequently, during backpropagation, during backpropagation, and can be optimized as follows:
where denotes the loss and denotes the learning rate. After fine-tuning, SparseLoRA first prunes with binary masks and then incorporate it with the pruned weights : . The adaptation of SparseLoRA finetuning ensures the sparsity of incremental weights, thus preserving the sparse pattern after merging. Other than the vision model and language model, VLMs also involve small learnable interfaces (e.g., QFormer Li et al. (2023); Dai et al. (2023)) that align vision models and language models. Because of this, we also insert LoRA into the QFormer, which enhances cross-modality adaptation with minimal additional computational overhead.
4.3 Cross-modality Adaptation
To recover the performance of pruned VLMs, we introduce two finetuning objectives. Firstly, acknowledging the performance gap, we continue to finetune VLMs on the pretraining tasks by minimizing loss to restore task-specific performance. On the other hand, we propose distilling knowledge Hinton et al. (2015); Gou et al. (2021); Stanton et al. (2021) from the original models to the pruned models by constraining the KL divergence between their outputs. The distillation loss is formulated as follows:
where represents the KL-divergence distance and logits denotes the output logits of the model with weights . Based on the original model with weights , both and can be obtained by forwarding with and separately. This avoids hosting additional weights during training. The overall optimization objective of RESSA is:
where is a scalar weight. The overall procedure of RESSA is shown in Figure 4.
5 Experimental Setup
Architectures. We use multiple multi-modal architectures for experiments including BLIP-2 Li et al. (2023) and InstructBLIP Dai et al. (2023), which composes of pretrained EVA-ViT (ViT-g/14 from EVA-CLIP) Sun et al. (2023b) and pretrained large language models (i.e., FlanT5 Chung et al. (2022) and Vicuna Chiang et al. (2023)).
Evaluation Datasets and Metrics. We evaluate the zero-shot ability of BLIP-2 and InstructBLIP on various datasets after pruning. We use VQAv2 Goyal et al. (2016), OK-VQA Marino et al. (2019), and GQA Hudson & Manning (2019) for visual question answering, NoCaps Agrawal et al. (2019) for image captioning, and Flickr30k Plummer et al. (2015) for image-text retrieval. We use CIDEr and SPICE to evaluate image captioning tasks and use TR@1 (top-1 text recall) and IR@1 (top-1 image recall) for image retrieval tasks.
Calibration and Training Datasets. Following Yi-Lin Sung (2024); Liu et al. (2023), our approach leverages a small subset of CC3M Sharma et al. (2018) for calibration and training data. The number of training samples ranges from 1k to 10k, while the number of calibration samples is 128, which has been shown to be sufficient for pruning Sun et al. (2023a); Frantar & Alistarh (2023); Zhang et al. (2024); Yi-Lin Sung (2024).
Finetuning Details. We use Adam Kingma & Ba (2015) as the optimizer with , = 0.9, 0.999. For regularization, we set the as 0.1 and grid-search the learning rate from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4}, where we warm up the learning rate in the first 10% steps (of the total training steps). For different model scales, we select a batch size from {16, 32, 64}, and finetune 1 epoch, which is enough for convergence. We perform a grid search for the rank of SparseLoRA, considering values from {4, 8, 16, 32}. By trial and error, we found that a rank of 4 suffices for the QFormer and the vision model, while a rank of 8 optimally suits the language model.
[…]
6 Results
| Method | Sparsity | Visual Question Answering | Image Captioning | Image retrieval | Macro Avg. | ||||
| VQAv2 | OK-VQA | GQA | NoCaps | Flickr30k | |||||
| Accuracy | CIDEr | SPICE | TR@1 | IR@1 | |||||
| Full Model | 0% | 63.1 | 41.1 | 44.1 | 105.4 | 13.8 | 96.1 | 87.5 | 64.4 |
| Magnitude | 50% | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.2 | 0.1 | 0.0 |
| Gradient | 55.1 | 35.7 | 39.8 | 92.3 | 11.6 | 91.4 | 81.6 | 58.2 | |
| SparseGPT | 56.1 | 35.5 | 40.6 | 98.7 | 13.3 | 95.8 | 86.2 | 60.9 | |
| Wanda | 57.7 | 35.4 | 41.9 | 100.1 | 13.4 | 95.2 | 84.5 | 61.2 | |
| ECoFLaP | 57.5 | 36.2 | 42.1 | 99.0 | 12.5 | 95.7 | 85.8 | 61.3 | |
[…]
| RESSA | 61.2 | 39.5 | 43.5 | 106.6 | 14.1 | 96.0 | 87.2 | 64.0 | |
6.1 Main Experimental Results
Unstructured Sparsity. In Table 1, We compare the zero-shot performance on various datasets using BLIP-2 pruned by different pruning techniques at unstructured 50% sparsity ratios. Among all pruning methods, while Wanda and ECoFLaP achieve the best performance, Wanda does not require multiple forward passes and is much more time-efficient. On the other hand, considering, EcoFLaP does not apply for N:M sparsity, we use Wanda as the default pruning method.
[…]
| Method | VQA | NoCaps | ||
| VQAv2 | GQA | CIDEr | SPICE | |
| Full Model | 76.7 | 49.1 | 123.9 | 15.9 |
| 2:4 Sparsity | ||||
[…]
| RESSA | 68.3 | 45.4 | 119.3 | 15.5 |
| 4:8 Sparsity | ||||
[…]
| RESSA | 71.4 | 46.5 | 121.6 | 15.6 |
N:M Sparsity. In addition to unstructured sparsity, we also conduct experiments on N: M sparsity andYukun Ma et al. (2021); Zhang et al. (2022), which can be applied to specific GPU cores and has more practical applications Mishra et al. (2021). Compared to unstructured pruning, structured pruning causes a more significant performance drop and requires more extensive repairs. Under more structured patterns, RESSA recovers more performance, achieving a 10.5% improvement for 2:4 sparsity compared to 3.4% for unstructured sparsity. After repairing, all structured pruned models maintain over 90% of the performance of the original models, demonstrating the universality and effectiveness of RESSA.
Pruning Single Modality. Language models typically have much larger parameter sizes compared to the vision models in vision-language models, Li et al. (2023); Dai et al. (2023) (e.g., 7B for Vicuna Chiang et al. (2023) vs. 1.3B for EVA-ViT in parameters Sun et al. (2023b)). As a result, the efficiency bottleneck primarily stems from the language model component. This prompted us to investigate the impact of solely pruning language models in VLMs, with experimental results presented in Table 2. With additional parameters in the vision model component, RESSA repairs InstructBLIP with significant improvement (e.g., from 69.0 to 71.6 on VQAv2), achieving performance comparable to the Full Model. Therefore, pruning language models only is an effective way to maintain performance and efficiency.
| Method | Sparsity | Visual Question Answering | Image Captioning | Macro Avg. | ||||
| VQAv2 | OK-VQA | GQA | NoCaps | |||||
| Accuracy | CIDEr | SPICE | ||||||
| Full Model | 0% | 73.5 | 52.6 | 48.4 | 121.4 | 15.6 | 62.3 | |
| Wanda | 50% | 69.1 | 45.4 | 45.7 | 108.7 | 14.2 | 56.6 | |
[…]
| InstructBLIP | RESSA | |||||||
| Wanda | 2:4 | 61.2 | 33.9 | 42.1 | 82.5 | 11.9 | 46.3 | |
[…]
| RESSA | ||||||||
| Wanda | 4:8 | 66.0 | 39.8 | 45.1 | 97.1 | 13.1 | 52.2 | |
[…]
| RESSA | ||||||||
| Full Model | 0% | 76.7 | 58.8 | 49.1 | 123.9 | 15.9 | 64.9 | |
| Wanda | 50% | 67.7 | 47.8 | 44.9 | 109.7 | 14.6 | 56.9 | |
[…]
| InstructBLIP Vicuna-7B | RESSA | |||||||
| Wanda | 2:4 | 58.7 | 32.1 | 39.0 | 68.8 | 12.9 | 42.3 | |
[…]
| RESSA | ||||||||
| Wanda | 4:8 | 61.4 | 39.5 | 42.4 | 95.5 | 13.6 | 50.5 | |
[…]
| RESSA | ||||||||
6.2 Ablation Studies
| Method | Sparsity | VQAv2 | OK-VQA | GQA |
| Wanda | 4:8 | 54.4 | 39.5 | 42.4 |
[…]
| RESSA-L | 66.4 | 46.7 | 44.3 | |
| RESSA-Q | 62.5 | 40.2 | 43.3 | |
[…]
| RESSA-VLQ | 69.5 | 47.4 | 45.8 |
Cross-Modality Adaptation. To assess the efficacy of cross-modality adaptation, we integrate SparseLoRA into various models. Specifically, we denote the QFormer, vision model, and language model as ”Q”, ”V” and ”L” respectively. Different insertion settings are represented using combinations of these notations (e.g., “QLV” and “LV”). As shown in Table 4, finetuning within a single model contributes to repairing the performance, and finetuning models across different modalities further enhances the performance. Moreover, adjusting the QFormer, which aligns the dense vision model and language model, improves the cross-modality ability of sparse VLMs. In summary, jointly finetuning modality-specific models and the learnable interface contribute most to performance repair.
| Method | Sparsity | VQAv2 | OK-VQA | GQA |
| Full Model | 0% | 76.7 | 58.8 | 49.1 |
[…]
| SparseLoRA | 50% | 74.0 | 53.3 | 48.6 |
[…]
| SparseLoRA | 2:4 | 68.3 | 44.6 | 45.4 |
[…]
| SparseLoRA | 4:8 | 71.4 | 49.2 | 46.5 |
SparseLoRA Finetuning. LoRA weights cannot be merged with pruned weights, as this would disrupt the sparse pattern. Consequently, the presence of remaining LoRA modules leads to latency and slows down inference significantly Dery et al. (2024); Rücklé et al. (2021). To address this issue, SparseLoRA aims to resolve the unmerged weights of LoRA and eliminate the latency caused by LoRA modules. Table 5 compares the performance of LoRA with SparseLoRA for VLMs with sparse language models. Remarkably, SparseLoRA finetuning achieves improved performance with fewer trainable parameters, consistent with findings from He et al. (2022).
| Method | Flickr30k | NoCaps | ||
| TR@1 | IR@1 | CIDEr | SPICE | |
| Full Model | 96.1 | 87.5 | 105.4 | 13.8 |
[…]
| 95.4 | 86.6 | 102.2 | 13.4 | |
| 96.0 | 87.2 | 106.6 | 14.1 | |
Finetuning Objectives. We further investigate the impact of the proposed finetuning objectives on BLIP-2-. In Table 6, we consider three finetuning objectives: , , and . guides the task-specific performance while guides knowledge transferring from the original full model to the pruned dense model. either minimizing or improves the performance. In addition, jointly minimizing and helps the pruned models further recover performance.
Sparsity. To assess the effectiveness of RESSA across a broader range of sparsity ratios, we experimented on InstructBLIP-Vicuna-7B with unstructured sparsity ratios ranging from 40% to 80%. As the sparse ratio exceeded 50%, the performance of pruned models began to deteriorate, eventually collapsing when , highlighting the necessity of repairing. In such scenarios, RESSA significantly improved performance, particularly for higher sparsity ratios, achieving a recovery of 47.61% of scores at and 32.65% at .
Training samples. RESSA utilizes calibration datasets for pruning and retraining. We conducted experiments to explore the impact of the number of training samples. Specifically, we randomly sampled ( = 0, 100, 1k, 10k, 100k) training data points from CC3M Sharma et al. (2018) to finetune with 50% sparsity and report the average performance of visual question answering and image caption. As shown in Figure 7, we found that finetuning pruned VLMs with few-shot samples (i.e., 100) can improve performance by a substantial margin. Further finetuning with 10k training data points resulted in a significant boost in cross-modality ability. This suggests that a small amount of data is sufficient to repair the pruned vision-language models, leveraging the knowledge and capabilities acquired during pretraining Zhou et al. (2023). When 10k, the model’s capability continues to improve with more training data and gradually becomes saturated.
7 Conclusion
In this paper, motivated by the challenges associated with deploying VLMs in real-world applications, we investigate the potential of pruning VLMs. Specifically, recognizing that VLMs encompass models from different modalities, we conduct empirical studies to explore the distribution of sparsity ratios across these models and how sparsity impacts performance, thereby highlighting the necessity of repairing pruned VLMs. Subsequently, we introduce RESSA, which addresses this challenge by repairing pruned VLMs through cross-modality adaptation and SparseLoRA finetuning. Extensive experiments validate the effectiveness of RESSA, providing valuable insights for future research on VLM sparsity.

### v2

@@
-RESSA: Repair Sparse Vision-Language Models via Sparse Cross-Modality Adaptation
+Rethinking Pruning for Vision-Language Models: Strategies for Effective Sparsity and Performance Restoration
 Abstract
-Vision-Language Models (VLMs), integrating diverse information from multiple modalities, have shown remarkable success across various tasks.
-However, deploying VLMs, comprising large-scale vision and language models poses challenges in resource-constrained scenarios.
-While pruning followed by finetuning offers a potential solution to maintain performance with smaller model sizes, its application to VLMs remains relatively unexplored, presenting two main questions: how to distribute sparsity across different modality-specific models, and how to repair the performance of pruned sparse VLMs.
-To answer the first question, we conducted preliminary studies on VLM pruning and found that pruning vision models and language models with the same sparsity ratios contribute to nearly optimal performance.
-For the second question, unlike finetuning unimodal sparse models, sparse VLMs involve cross-modality interactions, requiring specialized techniques for post-pruning performance repair.
-Moreover, while parameter-efficient LoRA finetuning has been proposed to repair the performance of sparse models, a significant challenge of weights merging arises due to the incompatibility of dense LoRA modules with sparse models that destroy the sparsity of pruned models.
-To tackle these challenges, we propose to Repair Sparse Vision-Language Models via Sparse Cross-modality Adaptation (RESSA).
-RESSA utilizes cross-modality finetuning to enhance task-specific performance and facilitate knowledge distillation from original dense models.
-Additionally, we introduce SparseLoRA, which applies sparsity directly to LoRA weights, enabling seamless integration with sparse models.
-Our experimental results validate the effectiveness of RESSA, showcasing significant enhancements, such as an 11.3% improvement under 2:4 sparsity and a remarkable 47.6% enhancement under unstructured 70% sparsity.
-Code will be released at: https://github.com/Shwai-He/RESSA.
+Vision-Language Models (VLMs) integrate information from multiple modalities and have shown remarkable success across various tasks.
+However, deploying large-scale VLMs in resource-constrained scenarios is challenging.
+Pruning followed by finetuning offers a potential solution but remains underexplored for VLMs.
+This study addresses two key questions: how to distribute sparsity across different modality-specific models, and how to restore the performance of pruned sparse VLMs.
+Our preliminary studies identified two effective pruning settings: applying the same sparsity to both vision and language models, and pruning only the language models.
+While LoRA finetuning aims to restore sparse models, it faces challenges due to incompatibility with sparse models, disrupting the pruned sparsity.
+To overcome these issues, we propose SparseLoRA, which applies sparsity directly to LoRA weights.
+Our experimental results demonstrate significant improvements, including an 11.3% boost under 2:4 sparsity and a 47.6% enhancement under unstructured 70% sparsity.
+Code is released at: https://github.com/Shwai-He/VLM-Compression.
+Rethinking Pruning for Vision-Language Models: Strategies for Effective Sparsity and Performance Restoration
+Shwai He1 Ang Li1 Tianlong Chen2††thanks: Corresponding author 1University of Maryland, College Park 2University of North Carolina at Chapel Hill shwaihe@umd.edu, angliece@umd.edu, tianlong@cs.unc.edu
 1 Introduction
@@
 (2020), which reduces model size while preserving performance, holds promise for improving the real-world deployment of VLMs.
-While pruning followed by finetuning has significantly improved the efficiency of vision models Frankle & Carbin (2019); Kusupati et al.
+While pruning followed by finetuning has significantly improved the efficiency of vision models Frankle and Carbin (2019); Kusupati et al.
 (2020); Lee et al.
@@
 (2020); Sun et al.
-(2023a); Frantar & Alistarh (2023), the realm of Vision-Language Models (VLMs) remains relatively unexplored in terms of model pruning, prompting following questions: how to distribute sparsity ratios between different modality-specific models and how to repair the performance of prune sparse VLMs.
-For the first question, we conducted empirical studies focusing on pruning modality-specific models.
-Specifically, we maintained a constant summation of sparsity ratios for language models and vision models and then experimented with various combinations of sparsity ratios.
-Surprisingly, we discovered that pruning the vision model and language model with the same sparsity ratios contributes to nearly optimal performance.
-However, even so, pruning both vision models and language models simultaneously would significantly degrade performance, especially for structured sparsity patterns (e.g., N:M sparsity Zhang et al.
+(2023a); Frantar and Alistarh (2023), the realm of Vision-Language Models (VLMs) remains relatively unexplored in terms of model pruning, prompting following questions: how to distribute sparsity ratios between different modality-specific models and how to restore the performance of prune sparse VLMs.
+For the first question, we conducted empirical studies on pruning modality-specific models, experimenting with various combinations of sparsity ratios.
+Surprisingly, we found that applying the same sparsity ratios to both vision and language models yields nearly optimal performance.
+On the other hand, since language models are usually much larger than vision models, pruning only the language models offers a beneficial trade-off between performance and efficiency.
+However, as sparsity ratios increase, pruning significantly degrades performance, especially with structured sparsity patterns (e.g., N: M sparsity Zhang et al.
 (2022); andYukun Ma et al.
-(2021)).
-While this degradation can be mitigated by only pruning language models, there still exists a significant performance gap compared to the full model, demonstrating the importance of post-pruning repairing.
-Unlike finetuning unimodal sparse models, sparse VLMs involve cross-modality interactions, requiring specialized techniques for post-pruning performance repair.
-Moreover, while parameter-efficient LoRA finetuning has been proposed to repair the performance of sparse models, a significant challenge arises due to the incompatibility of dense LoRA modules with sparse models: merging LoRA modules with sparse models would destroy the sparse pattern while maintaining LoRA modules would cause extra latency and slow down the inference speed.
-In response to the above challenges, we propose Repair Sparse Vision-Language Models via Sparse Cross-Modality Adaptation (RESSA).
-Specifically, we define two repairing objectives: enhancing task-specific performance and minimizing divergence from the original dense model.
-The former objective guides the model to improve performance, while the latter facilitates knowledge transfer from dense models to pruned models.
+(2021)), underscoring the importance of post-pruning restoration.
+While parameter-efficient LoRA finetuning has been proposed to repair the performance of sparse models, it faces a significant challenge due to the incompatibility of dense LoRA modules with sparse models.
+Merging LoRA modules with sparse models would destroy the sparse pattern, while maintaining LoRA modules would introduce extra latency and slow down the inference speed.
 To address the incompatibility issue of LoRA, we introduce SparseLoRA finetuning, which utilizes binary masks on LoRA weights, allowing seamless integration with pruned weights.
-SparseLoRA is incorporated into both the vision and language model to ensure cross-modality finetuning.
-Extensive experiments showcase the effectiveness of our proposed RESSA in repairing the performance of pruned sparse VLMs.
-For instance, as illustrated in Figure 1, RESSA boosts the performance by 13.1% for InstructBLIP-Vicuana-7B with 2:4 sparsity.
+Extensive experiments showcase the effectiveness of our proposed methods in repairing the performance of pruned sparse VLMs.
+For instance, as illustrated in Figure 1, SparseLoRA boosts the performance by 13.1% for InstructBLIP-Vicuana-7B with 2:4 sparsity.
 In summary, our contributions are threefold:
 We empirically study the modality-specific sparsity distributions and systematically demonstrate how sparsity affects the performance of VLMs.
-Following the empirical studies, we propose RESSA that repairs the pruned models via cross-modality adaptation and SparseLoRA finetuning.
-Extensive experiments demonstrate the effectiveness and universality of RESSA across various VLMs and tasks.
+We propose a pipeline involving pruning and post-finetuning with SparseLoRA to restore pruned models.
+Extensive experiments validate the effectiveness and universality of SparseLoRA across various VLMs and tasks.
 2 Related Work
@@
 To mitigate this, model pruning techniques have been introduced to remove redundant weights or structures Han et al.
-(2016); Alvarez & Salzmann (2016).
+(2016); Alvarez and Salzmann (2016).
 The primary aim of model pruning is to minimize the disparity between models before and after pruning Liu et al.
-(2021); He & Xiao (2024); Frantar & Alistarh (2023).
+(2021); He and Xiao (2024); Frantar and Alistarh (2023).
 Various metrics, such as magnitude, gradient Yi-Lin Sung (2024), and activation Sun et al.
@@
 (2022) (i.e., in each contiguous block of values, values must be zero), all models encounter significant performance degradation and even collapse (2:4 for pruning both vision models and language models).
-This situation prompts us to contemplate the question of how to repair the performance degradation caused by pruning for VLMs.
+This situation prompts us to reflect on how to restore the pruning-caused performance degradation for VLMs.
 4 Methodology
-In this section, we will develop a pruning pipeline, RESSA, that repairs the pruned models via cross modality adaptation and SparseLoRA finetuning, with the illustration in Figure 4.
+In this section, we will develop a pipeline that involves pruning and post restoration, with the illustration in Figure 4.
 4.1 Pruning with Few Samples
@@
 (2016), typically measuring the reconstruction errors Sun et al.
-(2023a); Frantar & Alistarh (2023); Zhang et al.
+(2023a); Frantar and Alistarh (2023); Zhang et al.
 (2024) between models before and after pruning, such as magnitude, gradient, and activation.
@@
 Because of this, we also insert LoRA into the QFormer, which enhances cross-modality adaptation with minimal additional computational overhead.
-4.3 Cross-modality Adaptation
+4.3 Finetuning Objectives
 To recover the performance of pruned VLMs, we introduce two finetuning objectives.
@@
 This avoids hosting additional weights during training.
-The overall optimization objective of RESSA is:
-where is a scalar weight. The overall procedure of RESSA is shown in Figure 4.
+The overall optimization objective of SparseLoRA is:
+where is a scalar weight. The procedure of VLM pruning and SparseLoRA is shown in Figure 4.
+| Method | Sparsity | Param. | Visual Question Answering | Image Captioning | Image retrieval | Macro Avg. | ||||
+| VQAv2 | OK-VQA | GQA | NoCaps | Flickr30k | ||||||
+| Accuracy | CIDEr | SPICE | TR@1 | IR@1 | ||||||
+| Full Model | 0% | 3.9B | 63.1 | 41.1 | 44.1 | 105.4 | 13.8 | 96.1 | 87.5 | 64.4 |
+| Magnitude | 50% | 2.1B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.2 | 0.1 | 0.0 |
+| Gradient | 55.1 | 35.7 | 39.8 | 92.3 | 11.6 | 91.4 | 81.6 | 58.2 | ||
+| SparseGPT | 56.1 | 35.5 | 40.6 | 98.7 | 13.3 | 95.8 | 86.2 | 60.9 | ||
+| Wanda | 57.7 | 35.4 | 41.9 | 100.1 | 13.4 | 95.2 | 84.5 | 61.2 | ||
+| ECoFLaP | 57.5 | 36.2 | 42.1 | 99.0 | 12.5 | 95.7 | 85.8 | 61.3 | ||
+[…]
+| Wanda + SparseLoRA | B | 61.2 | 39.5 | 43.5 | 106.6 | 14.1 | 96.0 | 87.2 | 64.0 | |
 5 Experimental Setup
@@
 (2016), OK-VQA Marino et al.
-(2019), and GQA Hudson & Manning (2019) for visual question answering, NoCaps Agrawal et al.
+(2019), and GQA Hudson and Manning (2019) for visual question answering, NoCaps Agrawal et al.
 (2019) for image captioning, and Flickr30k Plummer et al.
@@
 The number of training samples ranges from 1k to 10k, while the number of calibration samples is 128, which has been shown to be sufficient for pruning Sun et al.
-(2023a); Frantar & Alistarh (2023); Zhang et al.
+(2023a); Frantar and Alistarh (2023); Zhang et al.
 (2024); Yi-Lin Sung (2024).
-Finetuning Details.
-We use Adam Kingma & Ba (2015) as the optimizer with , = 0.9, 0.999.
+Finetuning Details We use Adam Kingma and Ba (2015) as the optimizer with , = 0.9, 0.999.
 For regularization, we set the as 0.1 and grid-search the learning rate from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4}, where we warm up the learning rate in the first 10% steps (of the total training steps).
@@
 6 Results
-| Method | Sparsity | Visual Question Answering | Image Captioning | Image retrieval | Macro Avg. | ||||
-| VQAv2 | OK-VQA | GQA | NoCaps | Flickr30k | |||||
-| Accuracy | CIDEr | SPICE | TR@1 | IR@1 | |||||
-| Full Model | 0% | 63.1 | 41.1 | 44.1 | 105.4 | 13.8 | 96.1 | 87.5 | 64.4 |
-| Magnitude | 50% | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.2 | 0.1 | 0.0 |
-| Gradient | 55.1 | 35.7 | 39.8 | 92.3 | 11.6 | 91.4 | 81.6 | 58.2 | |
-| SparseGPT | 56.1 | 35.5 | 40.6 | 98.7 | 13.3 | 95.8 | 86.2 | 60.9 | |
-| Wanda | 57.7 | 35.4 | 41.9 | 100.1 | 13.4 | 95.2 | 84.5 | 61.2 | |
-| ECoFLaP | 57.5 | 36.2 | 42.1 | 99.0 | 12.5 | 95.7 | 85.8 | 61.3 | |
+| Method | Sparsity | Visual Question Answering | Image Captioning | Macro Avg. | ||||
+| VQAv2 | OK-VQA | GQA | NoCaps | |||||
+| Accuracy | CIDEr | SPICE | ||||||
+| Full Model | 0% | 73.5 | 52.6 | 48.4 | 121.4 | 15.6 | 62.3 | |
+| Wanda | 50% | 69.1 | 45.4 | 45.7 | 108.7 | 14.2 | 56.6 | |
 […]
-| RESSA | 61.2 | 39.5 | 43.5 | 106.6 | 14.1 | 96.0 | 87.2 | 64.0 | |
+| InstructBLIP | w/SparseLoRA | |||||||
+| Wanda | 2:4 | 61.2 | 33.9 | 42.1 | 82.5 | 11.9 | 46.3 | |
+[…]
+| w/SparseLoRA | ||||||||
+| Wanda | 4:8 | 66.0 | 39.8 | 45.1 | 97.1 | 13.1 | 52.2 | |
+[…]
+| w/SparseLoRA | ||||||||
+| Full Model | 0% | 76.7 | 58.8 | 49.1 | 123.9 | 15.9 | 64.9 | |
+| Wanda | 50% | 67.7 | 47.8 | 44.9 | 109.7 | 14.6 | 56.9 | |
+[…]
+| InstructBLIP Vicuna-7B | w/SparseLoRA | |||||||
+| Wanda | 2:4 | 58.7 | 32.1 | 39.0 | 68.8 | 12.9 | 42.3 | |
+[…]
+| w/SparseLoRA | ||||||||
+| Wanda | 4:8 | 61.4 | 39.5 | 42.4 | 95.5 | 13.6 | 50.5 | |
+[…]
+| w/SparseLoRA | ||||||||
 6.1 Main Experimental Results
@@
 […]
-| Method | VQA | NoCaps | ||
-| VQAv2 | GQA | CIDEr | SPICE | |
-| Full Model | 76.7 | 49.1 | 123.9 | 15.9 |
-| 2:4 Sparsity | ||||
+| Method | Param. | VQA | NoCaps | ||
+| VQAv2 | GQA | CIDEr | SPICE | ||
+| Full Model | 7.9B | 76.7 | 49.1 | 123.9 | 15.9 |
+| 2:4 Sparsity | |||||
 […]
 […]
-| RESSA | 68.3 | 45.4 | 119.3 | 15.5 |
-| 4:8 Sparsity | ||||
+| w/SparseLoRA | 68.3 | 45.4 | 119.3 | 15.5 | |
+| 4:8 Sparsity | |||||
 […]
 […]
-| RESSA | 71.4 | 46.5 | 121.6 | 15.6 |
+| w/SparseLoRA | 71.4 | 46.5 | 121.6 | 15.6 | |
 N:M Sparsity.
@@
 (2021).
-Compared to unstructured pruning, structured pruning causes a more significant performance drop and requires more extensive repairs.
-Under more structured patterns, RESSA recovers more performance, achieving a 10.5% improvement for 2:4 sparsity compared to 3.4% for unstructured sparsity.
-After repairing, all structured pruned models maintain over 90% of the performance of the original models, demonstrating the universality and effectiveness of RESSA.
-Pruning Single Modality.
+Compared to unstructured pruning, structured pruning causes a more significant performance drop and requires more extensive restores.
+Under more structured patterns, SparseLoRA recovers more performance, achieving a 10.5% improvement for 2:4 sparsity compared to 3.4% for unstructured sparsity.
+After restoring, all structured pruned models maintain over 90% of the performance of the original models, demonstrating the universality and effectiveness of SparseLoRA.
+Single Model Pruning.
 Language models typically have much larger parameter sizes compared to the vision models in vision-language models, Li et al.
@@
 This prompted us to investigate the impact of solely pruning language models in VLMs, with experimental results presented in Table 3.
-With additional parameters in the vision model component, RESSA repairs InstructBLIP with significant improvement (e.g., from 69.0 to 71.6 on VQAv2), achieving performance comparable to the Full Model.
+With additional parameters in the vision model component, SparseLoRA restores InstructBLIP with significant improvement (e.g., from 69.0 to 71.6 on VQAv2), achieving performance comparable to the Full Model.
 Therefore, pruning language models only is an effective way to maintain performance and efficiency.
-| Method | Sparsity | Visual Question Answering | Image Captioning | Macro Avg. | ||||
-| VQAv2 | OK-VQA | GQA | NoCaps | |||||
-| Accuracy | CIDEr | SPICE | ||||||
-| Full Model | 0% | 73.5 | 52.6 | 48.4 | 121.4 | 15.6 | 62.3 | |
-| Wanda | 50% | 69.1 | 45.4 | 45.7 | 108.7 | 14.2 | 56.6 | |
-[…]
-| InstructBLIP | RESSA | |||||||
-| Wanda | 2:4 | 61.2 | 33.9 | 42.1 | 82.5 | 11.9 | 46.3 | |
-[…]
-| RESSA | ||||||||
-| Wanda | 4:8 | 66.0 | 39.8 | 45.1 | 97.1 | 13.1 | 52.2 | |
-[…]
-| RESSA | ||||||||
-| Full Model | 0% | 76.7 | 58.8 | 49.1 | 123.9 | 15.9 | 64.9 | |
-| Wanda | 50% | 67.7 | 47.8 | 44.9 | 109.7 | 14.6 | 56.9 | |
-[…]
-| InstructBLIP Vicuna-7B | RESSA | |||||||
-| Wanda | 2:4 | 58.7 | 32.1 | 39.0 | 68.8 | 12.9 | 42.3 | |
-[…]
-| RESSA | ||||||||
-| Wanda | 4:8 | 61.4 | 39.5 | 42.4 | 95.5 | 13.6 | 50.5 | |
-[…]
-| RESSA | ||||||||
-6.2 Ablation Studies
-| Method | Sparsity | VQAv2 | OK-VQA | GQA |
-| Wanda | 4:8 | 54.4 | 39.5 | 42.4 |
-[…]
-| RESSA-L | 66.4 | 46.7 | 44.3 | |
-| RESSA-Q | 62.5 | 40.2 | 43.3 | |
-[…]
-| RESSA-VLQ | 69.5 | 47.4 | 45.8 |
-Cross-Modality Adaptation.
-To assess the efficacy of cross-modality adaptation, we integrate SparseLoRA into various models.
-Specifically, we denote the QFormer, vision model, and language model as ”Q”, ”V” and ”L” respectively.
-Different insertion settings are represented using combinations of these notations (e.g., “QLV” and “LV”).
-As shown in Table 4, finetuning within a single model contributes to repairing the performance, and finetuning models across different modalities further enhances the performance.
-Moreover, adjusting the QFormer, which aligns the dense vision model and language model, improves the cross-modality ability of sparse VLMs.
-In summary, jointly finetuning modality-specific models and the learnable interface contribute most to performance repair.
 | Method | Sparsity | VQAv2 | OK-VQA | GQA |
@@
 | SparseLoRA | 4:8 | 71.4 | 49.2 | 46.5 |
-SparseLoRA Finetuning.
+6.2 Detailed Analysis
+| Method | Modality | Vision + Language | Language | ||||||
+| VQAv2 | OK-VQA | GQA | Avg. | VQAv2 | OK-VQA | GQA | Avg. | ||
+| Wanda | – | 61.4 | 39.5 | 42.4 | 47.8 | 63.9 | 44.5 | 43.1 | 50.5 |
+| w/SparseLoRA | 64.3 | 42.6 | 45.8 | 51.0 | 66.0 | 44.7 | 43.3 | 51.3 | |
+| 66.4 | 46.7 | 44.3 | 52.5 | 70.8 | 46.5 | 49.5 | 55.6 | ||
+| 62.5 | 40.2 | 43.3 | 48.7 | 64.3 | 44.5 | 44.1 | 51.0 | ||
+| 69.5 | 47.4 | 45.8 | 54.2 | 70.6 | 46.3 | 49.3 | 55.4 | ||
+| 69.0 | 46.9 | 45.4 | 53.8 | 70.2 | 45.9 | 48.1 | 54.7 | ||
+To evaluate cross-modality adaptation, we integrate SparseLoRA into various models.
+Specifically, we denote the QFormer, vision model, and language model as "", "", and "" respectively.
+Different configurations are represented using combinations of these notations (e.g., "" and "").
+As shown in Table 5, for cross-modality pruning (i.e., Vision + Language), finetuning within a single model contributes to performance restoration, while finetuning models across two modalities further enhances performance.
+In cases of single modality pruning, finetuning the pruned model alone is sufficient for restoration.
+Notably, joint finetuning with the QFormer does not yield performance gains beyond finetuning the pruned models.
+| Method | Flickr30k | NoCaps | ||
+| TR@1 | IR@1 | CIDEr | SPICE | |
+| Full Model | 96.1 | 87.5 | 105.4 | 13.8 |
+[…]
+| 95.4 | 86.6 | 102.2 | 13.4 | |
+| 96.0 | 87.2 | 106.6 | 14.1 | |
+SparseLoRA Finetuning Achieves Comparable Performance with LoRA.
 LoRA weights cannot be merged with pruned weights, as this would disrupt the sparse pattern.
@@
 (2022).
-| Method | Flickr30k | NoCaps | ||
-| TR@1 | IR@1 | CIDEr | SPICE | |
-| Full Model | 96.1 | 87.5 | 105.4 | 13.8 |
-[…]
-| 95.4 | 86.6 | 102.2 | 13.4 | |
-| 96.0 | 87.2 | 106.6 | 14.1 | |
-Finetuning Objectives.
+The Effectiveness of Finetuning Objectives.
 We further investigate the impact of the proposed finetuning objectives on BLIP-2-.
@@
 In addition, jointly minimizing and helps the pruned models further recover performance.
-Sparsity.
-To assess the effectiveness of RESSA across a broader range of sparsity ratios, we experimented on InstructBLIP-Vicuna-7B with unstructured sparsity ratios ranging from 40% to 80%.
-As the sparse ratio exceeded 50%, the performance of pruned models began to deteriorate, eventually collapsing when , highlighting the necessity of repairing.
-In such scenarios, RESSA significantly improved performance, particularly for higher sparsity ratios, achieving a recovery of 47.61% of scores at and 32.65% at .
-Training samples.
-RESSA utilizes calibration datasets for pruning and retraining.
+Ablation Study on Sparsity.
+To assess the effectiveness of SparseLoRA across a broader range of sparsity ratios, we experimented on InstructBLIP-Vicuna-7B with unstructured sparsity ratios ranging from 40% to 80%.
+As the sparse ratio exceeded 50%, the performance of pruned models began to deteriorate, eventually collapsing when , highlighting the necessity of restoring.
+In such scenarios, SparseLoRA significantly improved performance, particularly for higher sparsity ratios, achieving a recovery of 47.6% of scores at and 32.7% at .
+Ablation Study on Calibration Datasets.
+SparseLoRA utilizes calibration datasets for retraining.
 We conducted experiments to explore the impact of the number of training samples.
@@
 Further finetuning with 10k training data points resulted in a significant boost in cross-modality ability.
-This suggests that a small amount of data is sufficient to repair the pruned vision-language models, leveraging the knowledge and capabilities acquired during pretraining Zhou et al.
+This suggests that a small amount of data is sufficient to restoration the pruned vision-language models, leveraging the knowledge and capabilities acquired during pretraining Zhou et al.
 (2023).
@@
 In this paper, motivated by the challenges associated with deploying VLMs in real-world applications, we investigate the potential of pruning VLMs.
-Specifically, recognizing that VLMs encompass models from different modalities, we conduct empirical studies to explore the distribution of sparsity ratios across these models and how sparsity impacts performance, thereby highlighting the necessity of repairing pruned VLMs.
-Subsequently, we introduce RESSA, which addresses this challenge by repairing pruned VLMs through cross-modality adaptation and SparseLoRA finetuning.
-Extensive experiments validate the effectiveness of RESSA, providing valuable insights for future research on VLM sparsity.
+Specifically, recognizing that VLMs encompass models from different modalities, we conduct empirical studies to explore the distribution of sparsity ratios across these models and how sparsity impacts performance, thereby highlighting the necessity of restoring pruned VLMs.
+Subsequently, we introduce MAF, which addresses this challenge by restoring pruned VLMs through cross-modality adaptation and SparseLoRA finetuning.
+Extensive experiments validate the effectiveness of MAF, providing valuable insights for future research on VLM sparsity.
+8 Limitations
+Despite our progress, limitations remain in our work.
+Although our proposed methods are universal for all VLM models, we have primarily focused on BLIP family models and selected tasks .
+We believe our methods can be easily extended to a broader range of models and tasks.
+On the other hand, given there may be potentially high-quality dataset for restoring pruned models, we believe the incorporation of such datasets would further promotes our proposed methods